A father is four time as old as his daughter. Five years ago he was seven times as old as his daughter. Find the present age of the father and daughter.

Answer by TimothyLamb(4379)

x = 4y
x - 5 = 7(y - 5)
---
put the system of linear equations into standard form
---
x = 4y
x - 4y = 0
x - 5 = 7y - 35
x - 7y = -30
---
x - 4y = 0
x - 7y = -30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x= 40
y= 10
---
answer:
father = 40
daughter = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php